# Two dimensional dimers beyond planarity

### v2

@@ 3.1 Definition
Graph. We assume that $r\in\mathbb{N}$. In order to define a generalized double dimer model corresponding to the spin system on the bipartite graph $G=(V,E)$, we need to enlarge the graph, by introducing a ghost vertex $g$, representing the effect of the external magnetic field, and a source vector, $s$, which allows us to represent correlation functions for the spin system in terms of dimers. Each such vertex is connected by an edge to a vertex of the original graph $G$. The enlarged graph is denoted by $G_{en}:=(V_{en},E_{en})$, with $V_{en}:=V\cup\{g,s\}$, $s\neq g$, $E_{en}:=E\cup E_{g}\cup E_{s}$, where $E_{g}:=\big{\{}\{x,g\}\,:\,x\in V\big{\}}$ and $E_{s}:=\big{\{}\{x,s\}\,:\,x\in V\big{\}}$. We write $x\stackrel{{\scriptstyle en}}{{\sim}}y$ for $\{x,y\}\in E_{en}$ and $x\sim y$ for $\{x,y\}\in E$. We recall that $V^{e}\subset V$ (resp. $V^{o}\subset V$) is the set of vertices having even (resp. odd) graph distance from the origin $o\in V$ in $G\subset G_{en}$.
Configurations. System realisations are triplets $w=(m^{1},m^{2},\pi)$, which are now defined. We first define the pairs $m=(m^{1},m^{2})\in\Sigma$, where
@@
$\Sigma:=\big{\{}(m^{1},m^{2})\in\mathbb{N}_{0}^{E_{en}}\times\mathbb{N}_{0}^{E_{en}}\,:\,\sum_{y\stackrel{{\scriptstyle en}}{{\sim}}x}m^{1}_{\{x,y\}}=\sum_{y\stackrel{{\scriptstyle en}}{{\sim}}x}m^{2}_{\{x,y\}}\,\,\mbox{ and }\,\,m_{\{x,g\}}^{i}\in r\mathbb{N}_{0}\,\,\,\forall x\in V,\forall i\in\{1,2\}\big{\}}.$
We interpret for each edge $e\in E_{en}$, $m^{i}_{e}$ as the number of dimers of colour $i\in\{1,2\}$ ($i$-dimers) on $e$. By definition, each vertex in $V$ can only be touched by the same number of blue and red dimers, such a constraint however does not hold for $s$ or $g$. Moreover, the number of dimers on each edge touching the ghost vertex is only allowed to be a multiple of $r$.
Given $m\in\Sigma$, we let $\mathcal{P}(m)$ be the set of pairing functions $\pi=(\pi_{x})_{x\in V}$ such that, at each odd vertex $x\in V^{o}$, $\pi_{x}$ is a bijection from the set of blue dimers touching $x$ to the set of red dimers touching $x$, and at each even vertex $x\in V^{e}$, $\pi_{x}$ is a bijection from the set of red dimers touching $x$ to the set of blue dimers touching $x$. Note that pairing functions are not associated to $g$ or $s$, hence the dimers touching these vertices are always ‘left unpaired’.
Hence, any triplet $w=(m^{1},m^{2},\pi)$ with $m=(m^{1},m^{2})\in\Sigma$ and $\pi\in\mathcal{P}(m)$ can be viewed as a collection of paths, where each path is a sequence of dimers with alternating colours. Each dimer in such a sequence is the image and it is mapped to a dimer of a different colour through the pairing function. Such a path can be ‘closed’ and lying completely in $G$ or ‘open’ and then start and end at a ghost or source vertex. The ghost and the source vertex can then be viewed as sources of open paths. We will refer to the closed paths as loop and to the open ones as walk.
We let $\mathcal{W}$ be the set of configurations $w=(m,\pi)=(m^{1},m^{2},\pi)$ such that $m=(m^{1},m^{2})\in\Sigma$ and $\pi\in\mathcal{P}(m)$.
@@
Local time and source vector. Given $m=(m^{1},m^{2})\in\Sigma$, we define the local time at each vertex $x\in V$ as
$n_{x}(m)=\sum_{\begin{subarray}{c}y\stackrel{{\scriptstyle e}}{{\sim}}x\end{subarray}}m^{1}_{\{x,y\}}=\sum_{\begin{subarray}{c}y\stackrel{{\scriptstyle e}}{{\sim}}x\end{subarray}}m^{2}_{\{x,y\}},$ (3.1)
corresponding to the total number of blue or red dimers touching that vertex. Given an element $m\in\Sigma$, we let $\partial m=\big{(}(\partial m)^{1}_{x},(\partial m)^{2}_{x}\big{)}_{x\in V}\in\mathbb{Z}^{V\times V}$ be the source vector, where
$(\partial m)^{i}_{x}:=m^{i}_{\{x,s\}}\in\mathbb{N}_{0}$
corresponds to the number of $i$-dimers on $\{x,s\}$. Note that these quantities are not defined for $x=g$ or $x=s$. By a slight abuse of notation sometimes we will use $n_{x}(w)$ and $(\partial w)^{i}_{x}$ rather than $n_{x}(m)$ and $(\partial m)^{i}_{x}$ for $w=(m,\pi)\in\mathcal{W}$, and we consider $n_{x}$ and $(\partial w)^{i}_{x}$ them as random variables in $\mathcal{W}$. We also use $m^{i}_{e}(w)$ for the random variable corresponding to the number of $i$ dimers on $e$.
@@
Measures and two-point functions. We define the (unnormalised, positive) measure in $\mathcal{W}$ assigning to each $w=(m,\pi)\in\mathcal{W}$ the weight
$\mu_{\beta,h,r}(w):=\tilde{\mu}_{\beta,h,r}(m)\prod_{x\in V}\frac{U\big{(}k_{x}(w)\big{)}}{n_{x}(w)!},$ (3.2)
where $k_{x}(w):=n_{x}(w)-m_{\{x,g\}}^{2}(w)$ for each $x\in V$, and
$\tilde{\mu}_{\beta,h,r}(m):=\prod_{e\in E}\frac{\beta^{m_{e}^{1}}}{m^{1}_{e}!}\frac{\beta^{m_{e}^{2}}}{m^{2}_{e}!}\prod_{x\in V}\frac{(\beta h)^{\frac{m_{\{x,g\}}^{1}}{r}}}{\left(\frac{m^{1}_{\{x,g\}}}{r}\right)!}\frac{(\beta h)^{\frac{m_{\{x,g\}}^{2}}{r}}}{\left(\frac{m^{2}_{\{x,g\}}}{r}\right)!}.$
@@
Note that the weight assigned by $\mu_{\beta,h,r}(w)$ does not depend on $\pi$, but only on the number of dimers on each edge and on their colour. We use $\mu_{\beta,h,r}$ also to denote expectations, hence, given $f:\mathcal{W}\mapsto\mathbb{R}$, we define
$\mu_{\beta,h,r}(f):=\sum\limits_{w\in\mathcal{W}}\mu_{\beta,h,r}(w)f(w).$
We now define the partition function of the random walk loop soup with no sources
$\mathcal{Z}_{G,\beta,h,r}:=\mu_{\beta,h,r}(\{\partial m=(0,0)\}),$ (3.3)
a probability measure on the space of paths with no sources $\mathcal{W}^{\ell}:=\{w\in\mathcal{W}\,:\,\partial w=(0,0)\}$,
$\forall w\in\mathcal{W}^{\ell}\quad\mathcal{P}_{G,\beta,h,r}(w):=\frac{\mu_{\beta,h,r}(w)}{\mathcal{Z}_{G,\beta,h,r}},$ (3.4)
the expectation with respect to (3.4), $\mathcal{E}_{G,\beta,h,r}$, and the two-point functions
$\mathcal{G}_{G,\beta}^{(1)}(x,y):=\frac{\mu_{G,\beta,0,0}(x\rightarrow y)}{\mathcal{Z}_{G,\beta,0,0}}\quad\quad\quad\mathcal{G}_{G,\beta}^{(2)}(x,y):=\mathcal{P}_{G,\beta,0,0}(x\leftrightarrow y),$ (3.5)
where $\{x\leftrightarrow y\}\subset\mathcal{W}^{\ell}$ is the event that there exists a loop visiting both $x$ and $y$ and $\{x\rightarrow y\}\subset\mathcal{W}$ is the event that there exists an walk ending with a dimer on $\{x,s\}$ and a dimer on $\{y,s\}$ and all the remaining paths are loops. Note that these quantities are defined under the assumption $h=0$, hence possible walks are not allowed to end at the ghost vertex.
@@ 3.2 Special cases
Some choices of the weight function lead to well known models in statistical mechanics, combinatorics, probability theory, and physics. We will explain these connections in the most interesting case $h=0$, in which case the measure (3.2) takes a particularly simple form, with no dimer touching the ghost vertex almost surely and open paths only terminating at the source vertex when present. We will keep assuming that $G=(V,E)$ is bipartite.
Monomer double dimer model.
@@
We obtain the monomer double dimer model when $U(0)=U(1)=1$ and $U(n)=0$ for all $n>1$. Under this choice only configurations such that each vertex is either touched by precisely one blue and one red dimer or by no dimer are allowed. We let $\tilde{\Sigma}:=\big{\{}(m^{1},m^{2})\in\{0,1\}^{E}\times\{0,1\}^{E}:\sum_{y\sim x}m_{\{x,y\}}^{1}=\sum_{y\sim x}m_{\{x,y\}}^{2}\in\{0,1\},\forall x\in V\big{\}}.$ The partition function (3.3) then takes the simple form
$\mathcal{Z}_{G,\beta,0,0}=\sum\limits_{(m^{1},m^{2})\in\tilde{\Sigma}}\prod_{e\in E}\beta^{m^{1}_{e}+m^{2}_{e}}=\beta^{|V|}\,\mathbb{Z}_{G,1,\frac{1}{\beta}},$
@@
where $\mathbb{Z}_{G,1,\frac{1}{\beta}}$ was defined in (1.1). For the first identity above we used that $h=0$ and the fact that, given a realisation $m\in\Sigma$ such that each vertex $x\in V$ is either touched by precisely one red and blue dimer or is left untouched, $\mathcal{P}(m)$ contains just one element. For the second identity we observed that the number of monomers plus the total number of dimers always equal $|V|$. Hence we deduce that the probability measure (3.4) then corresponds to the probability measure in $\tilde{\Sigma}$ assigning to each $m\in\tilde{\Sigma}$ the weight
$\frac{\prod_{e\in E}\beta^{m_{e}^{1}+m_{e}^{2}}}{\mathcal{Z}_{G,\beta,0,0}}.$
and that, for each pair of vertices $x,y\in V$,
$\mathcal{G}^{(2)}_{G,\beta}(x,y)=\mathbb{P}_{G,1,\frac{1}{\beta}}(x\leftrightarrow y),$ (3.6)
where we recall the definition (1.1) and, by a slight abuse of notation, we use $\{x\leftrightarrow y\}\subset\Omega$ also to denote the event that there exists a loop visiting both $x$ and $y$ in the monomer double dimer model.
Recalling the description of the monomer double dimer model as a dimer model on the duplicated torus, the two-point function $\mathcal{G}^{(1)}_{G,\beta}(x,y)$ can be viewed as the ratio between the weight of all dimer configurations on the duplicated torus with monomers displaced at two vertices of such a graph and the weight of all configurations with no such monomers.
Double dimer model and perfect matchings.
@@
We obtain the double dimer model (corresponding to the monomer double dimer model with zero monomer activity) if $U(1)=1$ and $U(n)=0$ for each $n\neq 0$, in other words each vertex is touched by precisely one blue and red dimer. The function $g(s)$ defining the corresponding spin measure (2.4) is then
$g(s)=\frac{1}{(2\pi)^{2}}e^{-i(s^{1}-s^{2})}.$
Taking $\beta=1$ we have that the partition function (3.3) takes the simple form
$\mathcal{Z}_{G,1,0,0}=|\,\mathcal{D}_{G}(\emptyset)|^{2}.$
We further have that
$\mu_{G,1,0,0}(x\rightarrow y)=|\,\mathcal{D}_{G}(\emptyset)|\,|\,\mathcal{D}_{G}(\{x,y\})|$
@@
(see also [30, Figure 2]). Hence from the definition (3.5) and the two previous expressions we obtain the following identity, which is key for the proof of Theorem 1.2
$\mathcal{G}_{G,1}^{(1)}(x,y)=\frac{|\,\mathcal{D}(\{x,y\})|}{|\,\mathcal{D}(\emptyset)|}$ (3.7)
and holds for any pair of distinct vertices $x,y\in V$.
@@
We obtain the double dimer representation of the XY model when $U(n)=1$ for each $n\in\mathbb{N}_{0}$. In this case the random loop model (2.4) is equivalent to the ones which have been defined in [12, 23] (to see this connection, one should interpret the two colours of the dimers as the two possible directions of the edges in the multigraph which is considered in [12]), where it is proved that such models are a representation of the XY model.

### v3

@@ 3.1 Definition
In order to express correlations of our general spin measure in terms of realisations of the multi-occupancy double-dimer model we need to enlarge the (bipartite) graph $G=(V,E)$.
Definition 3.1 (Enlarged graph).
We enlarge $G$ by adding two auxiliary vertices:
- a ghost vertex $g$, representing the effect of the external magnetic field;
- a source vertex $s$, which allows us to express correlation functions of the spin system in terms of dimers.
Each of these vertices is connected by an edge to every vertex of the original graph $G$. We denote the enlarged graph by $G_{en}=(V_{en},E_{en})$, where
$V_{en}:=V\cup\{g,s\},\qquad g\neq s,\hskip 18.49988ptE_{en}:=E\cup E_{g}\cup E_{s},$
with
$E_{g}:=\bigl\{\{x,g\}:x\in V\bigr\},\hskip 18.49988ptE_{s}:=\bigl\{\{x,s\}:x\in V\bigr\}.$
We will refer to $G=(V,E)$ as the original graph, in order to distinguish it from the enlarged graph.
3.1.1 Configurations
We now introduce the configurations of the generalised monomer double-dimer model, which correspond to the spin measure defined in Section 2. The definition of the configuration space depends on the parameter $r$ — which is fixed to be an integer— appearing in the spin measure (2.5).
Definition 3.2 (Dimer cardinalities).
We first introduce the set of dimer cardinalities
@@ 3.1.1 Configurations
$\Sigma_{r}:=\Bigl\{(m^{1},m^{2})\in\mathbb{N}_{0}^{E_{en}}\times\mathbb{N}_{0}^{E_{en}}\,:\,\sum_{y\stackrel{{\scriptstyle en}}{{\sim}}x}m^{1}_{\{x,y\}}=\sum_{y\stackrel{{\scriptstyle en}}{{\sim}}x}m^{2}_{\{x,y\}}\ \ \text{and}\ \ m^{i}_{\{x,g\}}\in r\mathbb{N}_{0},\ \ \forall x\in V,\ i\in\{1,2\}\Bigr\},$
where the sum is over all vertices $y$ which are neighbours of $x$ in the enlarged graph (hence, including the ghost and the source vertex).
For each edge $e\in E_{en}$, the value $m^{i}_{e}$ is interpreted as the number of dimers of colour $i\in\{1,2\}$ (ii-dimers) occupying $e$.
By construction, each vertex $x\in V$ must be incident to the same number of blue and red dimers. This constraint does not apply to $s$ or $g$, which are elements of $V^{en}$ but not of $V$. Moreover, the number of edges incident to $g$ must carry a multiple of $r$ dimers.
Definition 3.3 (Match functions).
Given $m\in\Sigma$, for each $x\in V$ we define $\mathcal{P}_{x}(m)$ as the set of match functions at $x$, namely the set bijections between the blue dimers incident to $x$ and the red dimers incident to $x$. We define the match functions as the Cartesian product $\mathcal{P}(m)=\prod_{x\in V}\mathcal{P}_{x}(m)$.
Match functions match at each vertex of the original graph one blue dimer with a corresponding red dimer. This allows us to interpret configurations a collections of open or closed paths.
Definition 3.4 (Configurations).
We denote by $\mathcal{W}$ the set of configurations
$\mathcal{W}:=\{\,w=(m,\pi)=(m^{1},m^{2},\pi):m\in\Sigma_{r},\ \pi\in\mathcal{P}(m)\,\}.$
Each path is obtained by successively pairing dimers of opposite colours. A path may be either closed, lying entirely within $G$, or open, in which case it starts and ends at $g$ or $s$. This is because no pairing functions are defined at $g$ and $s$, and the constraint requiring the number of blue dimers to equal the number of red dimers does not necessarily hold at these vertices. Consequently, the ghost and source vertices serve as sources of open paths. We refer to closed paths as loops, and to open ones as walks.
3.1.2 Measure on double dimers
We now introduce a measure on the space of configurations. As a preliminary step, we define the local time.
Definition 3.5 (Local time).
@@ 3.1.2 Measure on double dimers
Given $m=(m^{1},m^{2})\in\Sigma_{r}$, the local time at each vertex $x\in V$ of the original graph is
$n_{x}(m)\;=\;\sum_{y\stackrel{{\scriptstyle en}}{{\sim}}x}m^{1}_{\{x,y\}}\;=\;\sum_{y\stackrel{{\scriptstyle en}}{{\sim}}x}m^{2}_{\{x,y\}}.$ (3.1)
That is, $n_{x}(m)$ equals the total number of blue (or equivalently red) dimers incident to $x$, including those connected to the ghost or source vertex. By slight abuse of notation, for $w=(m,\pi)\in\mathcal{W}$ we also write $n_{x}(w)$ in place of $n_{x}(m)$.
We now introduce a measure on $\mathcal{W}$.
Definition 3.6 (Measure).
@@
Let $r\in\mathbb{N}$ and $h,\beta\geq 0$. We define the (unnormalised, non-negative) measure on $\mathcal{W}$ by assigning to each configuration $w=(m,\pi)\in\mathcal{W}$ the weight
$\mu_{\beta,h,r}(w)\;:=\;\nu_{\beta,h,r}(m)\prod_{x\in V}\frac{U\bigl(k_{x}(w)\bigr)}{n_{x}(w)!},$ (3.2)
where
$k_{x}(w):=n_{x}(w)-m^{2}_{\{x,g\}}(w),\hskip 18.49988ptx\in V,$
and
$\nu_{\beta,h,r}(m)\;:=\;\prod_{i=1}^{2}\Big(\prod_{e\in E}\frac{\beta^{m_{e}^{i}}}{m^{i}_{e}!}\prod_{x\in V}\frac{(\beta h)^{\,m^{i}_{\{x,g\}}/r}}{\bigl(m^{i}_{\{x,g\}}/r\bigr)!}\Big)$ (3.3)
The weight of the configurations factorises into two contributions: one depending only on edge cardinalities (eq. (3.3)), and one depending only on the local time vector. To obtain correspondence with the spin measure when $h>0$, one must subtract the number of red dimers incident to the ghost vertex when evaluating the weight function in (3.2), this requires the introduction of the function $k_{x}(w)$.
When $h=0$, no edge is allowed to touch the ghost vertex, and the second product in (3.3) equals $1$. In this case, the measure $\mu_{\beta,h,r}$ does not depend on $r$, and we therefore use the simplified notation $\mu_{\beta,0,0}$.
Since $m^{i}_{\{x,g\}}$ is a multiple of $r$, the ratios $m^{i}_{\{x,g\}}/r$ are necessarily integers.
@@
Note that the weight $\mu_{\beta,h,r}(w)$ does not depend on the specific matching $\pi$, but only on the number of dimers on each edge and their colour.
We use the same notation $\mu_{\beta,h,r}$ for expectations.
Definition 3.7 (Expectation).
For a function $f:\mathcal{W}\to\mathbb{R}$ we set
$\mu_{\beta,h,r}(f)\;:=\;\sum_{w\in\mathcal{W}}\mu_{\beta,h,r}(w)\,f(w).$
3.1.3 Two-point functions
As observed before, any configuration in $\mathcal{W}$ can be viewed as a collection of paths: either loops, entirely contained in $G$, or walks, which necessarily have $s$ and/or $g$ as endpoints. In order to represent correlation functions of the spin system introduced in Section 2, we need to impose further constraints involving the number of dimers on edges incident to the source vertex.
Definition 3.8 (Source vector).
Given $w=(m,\pi)\in\mathcal{W}$, we define the source vector
$\partial w=\bigl((\partial w)^{1}_{x},(\partial w)^{2}_{x}\bigr)_{x\in V},$
where $(\partial w)^{i}_{x}:\mathcal{W}\mapsto\mathbb{N}_{0}$ denotes the number of $i$-dimers on the edge $\{x,s\}$. In other words, for any configuration $w=(m^{1},m^{2},\pi)\in\mathcal{W}$ we have
$(\partial w)^{i}_{x}:=m^{i}_{\{x,s\}}.$
For example, take any configuration $w\in\mathcal{W}$ with no dimer touching the ghost vertex and with $(\partial w)^{1}=\delta_{x}+\delta_{y}$ and $(\partial w)^{2}=0$, where $x$ and $y$ are vertices of $V$ at odd distance. By definition of the source vector, this configuration has precisely one blue dimer on the edges $\{x,s\}$ and $\{y,s\}$, and no additional blue or red dimers incident to the source vertex. Since on the original graph the number of blue dimers incident to each vertex equals the number of red dimers and each dimer is matched, it follows that the configuration consists of a walk whose last steps lie on the edges $\{x,s\}$ and $\{y,s\}$, together with an arbitrary collection of loops entirely contained in $G$. If, on the other hand, $x$ and $y$ were at even distance, no such configuration could exist, since any walk would necessarily have end-dimers of different colours.
Definition 3.9 (Partition function and induced probability measure).
We denote by
$\mathcal{W}^{\ell}:=\{\,w\in\mathcal{W}:\partial w=(0,0)\,\}$
the set of source-free configurations. In the special case $h=0$, any configuration in this set containing a walk receives zero weight from (3.2). The partition function of the multi-occupancy double-dimer model is defined as the total weight of all source-free configurations, namely
${Z}^{path}_{G,\beta,h,r}:=\mu_{\beta,h,r}\bigl(\mathcal{W}^{\ell}\bigr).$ (3.4)
On $\mathcal{W}^{\ell}$ we define the probability measure
$P_{G,\beta,h,r}(w):=\frac{\mu_{\beta,h,r}(w)}{{Z}^{path}_{G,\beta,h,r}},\hskip 18.49988pt\forall\,w\in\mathcal{W}^{\ell},$ (3.5)
and denote expectation with respect to $P_{G,\beta,h,r}$ by $E_{G,\beta,h,r}$.
Our method allows us to obtain information on various two-point functions, but we will focus on the most important ones. The first two-point function is the ratio of the weight of configurations with a walk and the partition function. The second two-point function is the probability, with respect to the probability measure defined in Definition 3.9, that there exists a loop connecting two sites.
Definition 3.10 (Two-point functions).
For $x,y\in V$, the two-point functions are defined as
$\mathcal{G}^{(1)}_{G,\beta}(x,y):=\frac{\mu_{\beta,0,0}(\,x\to y\,)}{Z^{path}_{G,\beta,0,0}},\hskip 18.49988pt\mathcal{G}^{(2)}_{G,\beta}(x,y):=P_{G,\beta,0,0}(\,x\leftrightarrow y\,),$ (3.6)
where
  • $\{x\leftrightarrow y\}\subset\mathcal{W}^{\ell}$ is the event that there exists a loop visiting both $x$ and $y$;
  • $\{x\to y\}\subset\mathcal{W}$ is the event that there exists a walk with one endpoint given by a dimer on $\{x,s\}$ and the other by a dimer on $\{y,s\}$, while all remaining paths are loops entirely contained in the original graph $G$.
Note that these quantities are defined under the assumption $h=0$, so walks ending at the ghost vertex are excluded.
@@ 3.2 Special cases
Certain choices of the weight function give rise to well-known models in statistical mechanics, combinatorics, probability theory, and physics. We explain these connections in the most interesting case $h=0$, where the measure (3.2) takes a particularly simple form: no dimer touches the ghost vertex almost surely, and open paths (when present) terminate only at the source vertex.
Monomer double-dimer model.
@@
We obtain the monomer double-dimer model when $U(0)=U(1)=1$ and $U(n)=0$ for all $n>1$. Under this choice, only configurations in which each vertex is either incident to precisely one blue and one red dimer, or to no dimer at all, are allowed. We let
$\tilde{\Sigma}:=\Bigl\{(m^{1},m^{2})\in\{0,1\}^{E}\times\{0,1\}^{E}:\sum_{y\sim x}m_{\{x,y\}}^{1}=\sum_{y\sim x}m_{\{x,y\}}^{2}\in\{0,1\},\ \forall x\in V\Bigr\}.$
The partition function (3.4) then takes the simple form
$Z^{path}_{G,\beta,0,0}=\sum_{(m^{1},m^{2})\in\tilde{\Sigma}}\prod_{e\in E}\beta^{m^{1}_{e}+m^{2}_{e}}=\beta^{|V|}\,\mathbb{Z}_{G,1/\beta},$
@@
where $\mathbb{Z}_{G,1/\beta}$ was defined in (1.4). For the first identity we used $h=0$ and the fact that, given $m\in\Sigma_{r}$ with each vertex $x\in V$ either incident to one blue and one red dimer or left untouched, the set $\mathcal{P}(m)$ contains exactly one element. For the second identity we used that the total number of monomers plus the total number of dimers always equals $|V|$. Hence the probability measure (3.5) coincides with the probability measure on $\tilde{\Sigma}$ assigning to each $m\in\tilde{\Sigma}$ the weight
$\frac{\prod_{e\in E}\beta^{m_{e}^{1}+m_{e}^{2}}}{Z^{path}_{G,\beta,0,0}}.$
Moreover, for any $x,y\in V$ we have
$\mathcal{G}^{(1)}_{G,\beta}(x,y)$ $=\mathbb{G}_{G,1/\beta}(x,y)$ (3.7)
$\mathcal{G}^{(2)}_{G,\beta}(x,y)$ $=\mathbb{P}_{G,1/\beta}(x\leftrightarrow y).$ (3.8)
where, the quantities on the right-hand side involve the monomer double-dimer model and have been defined in Section 1.3.2.
Double-dimer model and dimer model.
@@
We obtain the double-dimer model (corresponding to the monomer double-dimer model with zero monomer activity) by setting $U(1)=1$ and $U(n)=0$ for all $n\neq 1$. In other words, each vertex is incident to precisely one blue dimer and one red dimer. For $\beta=1$ the partition function (3.4) reduces to
$Z^{path}_{G,1,0,0}=\bigl|\mathcal{D}_{G}(\emptyset)\bigr|^{2}.$ (3.9)
Moreover,
$\mu_{1,0,0}(x\to y)=\bigl|\mathcal{D}_{G}(\emptyset)\bigr|\,\bigl|\mathcal{D}_{G}(\{x,y\})\bigr|$
@@
(see also [50, Figure 2]). Therefore, from (3.6) and the previous identities we obtain the following relation, which is key for the proof of Theorem 1.1:
$\mathcal{G}^{(1)}_{G,1}(x,y)=\mathcal{C}_{G}(x,y)$ (3.10)
holding for any distinct $x,y\in V$, where on the right-hand side we have the monomer-monomer correlation for the dimer model, which has been defined in (1.1).
@@
We obtain the double-dimer representation of the XY model when $U(n)=1$ for all $n\in\mathbb{N}_{0}$. In this case the random loop model (2.5) is equivalent to the models introduced in [36, 33] (to see the correspondence, one interprets the two colours of dimers as the two possible directions of edges in the multigraph considered in [36]), where it is proved that such models provide a representation of the XY model.